OpenStudy (anonymous):

Solve to get x values: e^x=2e^(-x)+1

OpenStudy (anonymous):

e^x=2e^(-x)+1 \[ore e^{x}=\frac{ 2 }{e ^{x} }+1,put e ^{x}=y,\] solve for y and then for x.